3.00 L of water is heated to 96.00°C by the addition of 52.70kJ of heat. What was the initial heat of the temperature?
1
Expert's answer
2019-09-25T05:43:29-0400

From the heat absorbed equation

We have

"q=mc(T_2-T_1)"


"m=2000\ g\\c=4.184\ J\ g^{-1}\ \degree C^{-1}"

"T_2=369\ K\\q=52700\ \ J"

"T_2-T_1=\frac{q}{mc}=\frac{52700}{2000\times 4.184}=6.297\ K"


"T_1=T_2-6.297=369-6.297=362.7\ K"


Initial temprature"\ =362.7\ K"